OpenStudy (anonymous):

The area of a triangle inscribed in a circle is 60 cm^2 and the radius of the circumscribed circle is 8 cm. if the two sides of the inscribed triangle measure 10 cm and 12 cm, find the length of the third side. My answer is 12 cm? is it right? please help me thank you :)

OpenStudy (doc.brown):

A quick check. If triangle is 12 x 12 x 10 we can say the base is 10 and the height unknown. h^2 + 5^2 = 12^2 h^2 = 144 - 25 h= 10.9 a = bh = 5 x 10.9 = 54.5 That's not 60, try again.

OpenStudy (anonymous):

can you help me i can't really solve this :(

OpenStudy (anonymous):

i'll give a medal!! :))

OpenStudy (doc.brown):

Draw doesn't work.. so bare with me.

OpenStudy (doc.brown):

The 10 cm line touches the circle at two points, that means it makes a triangle of 8, 8, and 10. The second triangle is 8, 8, and 12. The third triangle is 8, 8, and x.